OpenStudy (anonymous):

help please! Evaluate the summation of 20 times 0.5 to the n minus 1 power, from n equals 3 to 12.. 9.99 19.95 29.98 39.96

OpenStudy (anonymous):

\[\sum_{n=3}^{12} 20 \times 0.5^{n-1}\] This is it, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@larryboxaplenty

OpenStudy (anonymous):

I can't help you right now, sorry, g2g for a sec someone else can help. I know the answer for forget the shortcut

OpenStudy (anonymous):

can someone please help me

OpenStudy (jhannybean):

\[\large 20 \cdot \left(0.5^{n-1}\right)\] start with plugging in \(n=3\) and work your way up to \(n=12\) by adding all the values.

OpenStudy (jhannybean):

\[\large 20\cdot (0.5^{3-1} + 0.5^{4-1}+0.5^{5-1}+...+0.5^{12-1})\]

OpenStudy (welshfella):

or you can use the formula for the sum of a geometric series Sn = a . (1 - r^n) ------ 1 - r

OpenStudy (welshfella):

here a = 20*0.5^2 , n = 11 and r = 0.5

OpenStudy (anonymous):

so would i get 9.99 as the answer

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

thank you

OpenStudy (welshfella):

yw